# Tricriticality and finite-size scaling in the triangular Blume-Capel ferromagnet

### v1

@@ I Introduction
$\mathcal{H}=-J\sum_{\langle xy\rangle}\sigma_{x}\sigma_{y}+\Delta\sum_{x}\sigma_{x}^{2}=E_{J}+\Delta E_{\Delta},$ (1)
where the spin variables $\sigma_{x}$ take on the values $-1,0$, or $+1$, $\langle xy\rangle$ indicates summation over nearest neighbors, and $J>0$ is the ferromagnetic exchange interaction. The parameter $\Delta$ is known as the crystal-field coupling and it controls the density of vacancies ($\sigma_{x}=0$). We should note here that the decomposition on the right-hand side of Eq. (1) into the bond-related and crystal-field-related energy contributions $E_{J}$ and $E_{\Delta}$, respectively, will turn out to be useful in the context of the numerical simulations discussed below. For $\Delta\rightarrow-\infty$, vacancies are suppressed and the model becomes equivalent to the simple Ising ferromagnet.
The phase boundary of the model in the crystal field–temperature ($\Delta,T$) plane separates the ferromagnetic from the paramagnetic phase. The ferromagnetic phase is characterized by an ordered alignment of ±$\pm 1$ spins. The paramagnetic phase, on the other hand, can be either a completely disordered arrangement at high temperature or a ±$\pm 1$-spin gas in a 00-spin dominated environment for low temperatures and high crystal fields. At high temperatures and low crystal fields, the ferromagnetic–paramagnetic transition is a continuous phase transition in the Ising universality class, whereas at low temperatures and high crystal fields, the transition is of first order [1, 2]. The model is thus a paradigmatic example of a system with a tricritical point $(\Delta_{\rm t},T_{\rm t})$ [3], where the two segments of the phase boundary meet. At zero temperature, it is clear that ferromagnetic order must prevail if its energy $zJ/2$ per spin (where $z$ is the coordination number) exceeds that of the penalty $\Delta$ for having all spins in the ±$\pm 1$ state. Hence the point $(\Delta_{0}=zJ/2,T=0)$ is on the phase boundary [2]. On the other hand, for zero crystal fields, the transition temperature is not exactly known, but it has been well-studied for a number of lattice geometries. For a picture of the phase diagram based on the data of the current work but also on some previous computations from Ref. [8], we refer the reader to Fig. 1, where we fix units by choosing $J=1$ and $k_{\rm B}=1$.
@@
Figure 1: Phase diagram of the two-dimensional spin-$1$ triangular Blume-Capel model in the plane of crystal-field coupling $\Delta$ and temperature $T$. The first-order transition line and tricritical point are determined from the field-mixing analysis using the Wang-Landau (WL) joint density of states. The inset magnifies the phase boundary near the tricritical point in the plane of $\beta=1/T$ and $\mu=\Delta/T$, estimating the field-mixing parameter $r$ by its slope. The critical points obtained from the multicanonical (MUCA) calculations in this work and the earlier Wang-Landau simulations [8] are also displayed in the second-order area, while the first-order transition points from the multicanonical calculations are not shown because they are visually identical to the Wang-Landau results. The complete list of the transition points is given in Table 1.
In this paper we are interested in the critical properties of the Blume-Capel model embedded in the triangular lattice, where $z=6$ and $\Delta_{0}=3$. For this particular case, Mahan and Girvin [35] were the first to apply position-space renormalization-group methods to solve the model. These authors estimated the critical frontier with the location of the tricritical point $(\Delta_{\rm t},T_{\rm t})=(2.686,1.493)$. Many years later, Du et al.[36] performed a further sophisticated analytical calculation using an expanded Bethe-Peierls approximation to find respectively the estimates $(\Delta_{\rm t},T_{\rm t})=(2.841,1.403)$ [37]. The refinement $(\Delta_{\rm t},T_{\rm t})=(2.925(8),1.025(10))$ was proposed in Ref. [8] via Monte Carlo simulations and an approximate scheme that takes advantage of the scaling behavior of the specific heat.
@@ II.1 Wang-Landau joint density of states
Figure 2 displays a typical landscape of $\Gamma(E_{J},E_{\Delta})$ for a triangular lattice with $L=18$, revealing a sharp edge with a significant magnitude of $\Gamma$ that was not observed in the case of the square-lattice model [20, 24]. The structure of $\Gamma(E_{J},E_{\Delta})$ appearing in the positive $E_{J}$ area reflects the frustration associated with antiferromagnetic coupling. Although we focus on the ferromagnetic model, this structure affects the Wang-Landau process since it needs to explore the entire energy space, slowing down the sampling dynamics needed to achieve the flatness criterion. Here, Wang-Landau data of the joint density of states were obtained for linear sizes $L=\{8,9,10,12,14,15,16,18,20,21,22,24\}$. Five (two) samples of the density of states were prepared from independent simulations for $L<20$ ($L\geq 20)$. The deviations between different samples turned out invisibly small for the data analyzed in the following sections. In computational time cost, the largest system size $L=24$ of our current simulations in the triangular model is comparable to the previous calculations conducted on the square-lattice model with $L=48$ [24].
Figure 2: The logarithm of the joint density of states, $\Gamma(E_{J},E_{\Delta})$, for the triangular Blume-Capel ferromagnet. The results are derived from Wang-Landau simulations for a system with linear size $L=18$.
@@ II.3 Method of field mixing
where $s$ and $r$ are the system size-dependent field-mixing parameters, and $\beta_{\rm t}$ and $\mu_{\rm t}$ indicate the tricritical point. The direction of the scaling field $g$ is tangent to the phase-coexistence curve. The field $\lambda$ is across the phase boundary but is not necessarily orthogonal to $g$. The scaling operators conjugate to $\lambda$ and $g$ are
@@
respectively, where $\epsilon\equiv L^{-d}E_{J}$ and $n\equiv L^{-d}E_{\Delta}$ with dimension $d=2$ in our model, satisfying the relation $\langle\mathcal{O}_{x}\rangle=-L^{-d}\partial_{x}\ln\mathcal{Z}$. The probability distribution of the operator $\mathcal{Q}$ is a convenient tool to determine the phase-coexistence curve and locate the tricritical point.
At a first-order transition point, the probability distribution $P(\mathcal{Q})$ conjugate to the field $\lambda$ in the direction across the phase boundary is double-peaked and symmetric. On the other hand, the distribution of the operator $\mathcal{E}$ conjugate to the field $g$ along the tangent of the phase boundary is bell-shaped. Although both follow the scaling Ansätze near the tricritical point, it was demonstrated that the symmetric double-peaked structure of $P(\mathcal{Q})$ works as a sensitive indicator of the phase coexistence when computed using the Wang-Landau joint density of states [24].
@@
Although the double-peaked structure is evident at temperatures both below and above the tricritical temperature $T_{\rm t}$, as illustrated for small systems in Fig. 3, the definite first- and second-order character of the transition can be undoubtedly distinguished by the finite-size behavior of $P(\tilde{\mathcal{Q}})$. At the first-order transition regime the peaks become sharper as $L$ increases (see also Figs. 9(a) and 9(b) below). In contrast, at the second-order transition, as shown in Fig. 3(d), the peaks become more broadened with increasing $L$, indicating that the double-peaked structure vanishes in the thermodynamic limit.
@@ II.5 Tricritical exponents
The identification of the tricritical point in our field-mixing analysis prompts us to verify the proposed tricritical Ising exponents. Near the tricritical point the probability distribution can be expressed as
@@
where $p^{\ast}_{L}$ is a universal scaling function, and $a$’s are non-universal factors (for further details on the field-mixing method we refer the interested reader to Refs. [42, 43, 44, 45]). The relevant eigenvalue exponents for the tricritical Ising universality class in two dimensions are known to be $y_{t}=9/5$, $y_{g}=4/5$, and $y_{h}=77/40$ [48, 49, 50, 51].
At the tricritical point, the finite-size scaling Ansätze for the variable $\mathcal{Q}$ can be expressed more explicitly as
@@
where $\mathcal{P}_{o}$ is scale-invariant. To examine this scaling form it is essential to have both mixing parameters, $r$ and $s$, to compute $\mathcal{Q}$. While we have obtained $s$ from the field-mixing analysis, the other parameter $r$ can be determined from the phase coexistence curve, as $-r$ corresponds to the slope at the tricritical point in the $\beta-\mu$ plane. The inset of Fig. 1 indicates $r=-3.2009(3)$. Figure 5(a) shows an excellent finite-size scaling collapse of $P_{L}(\mathcal{Q})$ with the given exponent $y_{t}=9/5$ at $T=0.985$, verifying the conjectured tricritical exponent in the triangular Blume-Capel ferromagnet.
@@ III.1 Multicanonical simulations
In our multicanonical simulations, we considered triangular lattices with periodic boundary conditions and linear sizes in the range $12\leq L\leq 128$. To thoroughly probe the characteristics of the transition on both sides of the phase boundary across the tricritical point, we performed simulations at various temperatures from $T=2.395$ in the second-order transition regime down to $T=0.75$ in the first-order regime.
We note here that the multicanonical simulations become prohibitively demanding in the low-temperature regime, a fact that is also reflected in the statistical fluctuations of the large-$L$ data illustrated in Figs. 10 and 11 for $T=0.75$. Last but not least, for all fits performed throughout this paper, we restricted ourselves to data with $L\geq L_{\rm min}$, adopting the standard $\chi^{2}$ test for goodness of the fit. Specifically, we considered a fit as being acceptable only if $10\%<Q<90\%$, where $Q$ is the quality-of-fit parameter [65].
@@ III.2 Observables
At this point, having defined some of the main observables of interest, we call attention to a comparative check that highlights the numerical accuracy of the simulation methods implemented in this work.
Figure 8 contrasts the specific heat and magnetic susceptibility curves of a $L=16$ triangular Blume-Capel system at the temperature $T=2.395$ obtained via three different protocols: (i) the Wang-Landau joint density of states outlined in Sec. II.1, (ii) the multicanonical scheme sketched here, and (iii) the well-known Metropolis algorithm [52]. Without doubt, Fig. 8 showcases a perfect matching among the three datasets.
As we are mostly interested in the first-order characteristics of the transition, following the prescription of Ref. [61], we constructed the probability density function of $E_{\Delta}$ per spin, $P(e_{\Delta})$, where $e_{\Delta}=E_{\Delta}/L^{2}$. It is well known that a double-peaked structure in the probability density function in finite systems is a precursor of the two delta functions expected in the thermodynamic limit for a first-order phase transition [46, 47]. Typical (normalized to unity) probability distributions $P(e_{\Delta})$ for various system sizes and temperatures in the first-order transition regime and at the proposed tricritical point are shown in Fig. 9. A few comments are in order: (i) In the first-order transition regime, the barrier increases upon lowering the temperature, reaching the order of $\sim 10^{-24}$ for $T=0.75$ [panel (a)]. (ii) Note the strong enhancement of the barrier with increasing system size [panel (b)], exhibiting a major suppression of intermediate states, which is common for first-order phase transitions. (iii) At the proposed tricritical point, [panel (c)], the distributions still feature a double-peaked structure, yet with a much smaller barrier which seems to be independent of system size. A more illuminating discussion on the asymptotic behavior of these observables regarding the distributions of Fig. 9 is given below.
@@ III.3.1 First-order transition regime
where $P_{\rm max}$ and $P_{\rm min}$ are the maximum and local minimum of the distribution $P(e_{\Delta})$, respectively. The resulting barrier connects a spin-00 dominated regime ($e_{\Delta}$ small) and a spin-±$\pm 1$ rich phase ($e_{\Delta}$ large). The corresponding surface tension $\Sigma(L)=\Delta F(L)/L$ is expected to scale as
@@
in two dimensions, possibly with higher-order corrections [72, 73, 74]. The scaling behavior of this observable is depicted in Fig. 10(a) for the complete set of temperatures considered during the multicanonical simulations and the solid lines are fits of the form (28). The retrieved $\Sigma_{\infty}$ values are plotted against the temperature difference from the tricritical point ($T_{\rm t}-T$) in Fig. 10(b), featuring an almost linear increase in $\Sigma_{\infty}$ as the temperature decreases towards the limit $T\rightarrow 0$.
This result agrees nicely with the observations by Jung and Kim for the square-lattice Blume-Capel model; see Fig. 4(a) in Ref. [33]. Additionally, we verify that at the proposed tricritical temperature $T=0.985$, the surface tension vanishes to a limiting value of $\Sigma_{\infty}=3\times 10^{-6}\pm 2\times 10^{-5}$.
We now examine the system-size scaling behavior of the pseudo-transition field $\Delta_{L}^{\ast}$ obtained using the probability density function $P(e_{\Delta})$. We determine $\Delta_{L}^{\ast}$ by the value of the crystal field at which the two peaks of equal height appear in $P(e_{\Delta})$, as illustrated in Fig. 9. Figure 11 presents the behavior of $\Delta_{L}^{\ast}$ for selected temperatures in the first-order transition regime. Our data indicate an excellent fit to the expected first-order behavior of the scaling form (10) as demonstrated by the solid lines in Fig. 11. The extrapolated values of the transition field $\Delta_{\infty}^{\ast}$ obtained here from $P(e_{\Delta})$ confirm to a high numerical accuracy the transition points obtained from the field-mixing analysis in Sec. II.3. Table 1 provides the comparison between the two different methods, showing that a deviation in $\Delta_{\infty}^{\ast}$ is about $10^{-4}$ or less.
@@
where $\{b_{C},b^{\prime}_{C},b_{\chi},b^{\prime}_{\chi}\}$ are non-universal amplitudes. This scaling Ansatz provides an excellent fit to our data for both observables, as shown in Fig. 12. From the data fitting with $L_{\mathrm{min}}=24$, we obtain $x=2.000(2)$ and $2.000(4)$ for the specific heat and the magnetic susceptibility, respectively, which are in perfect agreement with the expected value $d=2$ of the first-order transition in this model. Note that due to the value of $x\approx 2$, the $1/L^{2}$ correction simply becomes an additive constant.

### v2

@@ I Introduction
$\mathcal{H}=-J\sum_{\langle xy\rangle}\sigma_{x}\sigma_{y}+\Delta\sum_{x}\sigma_{x}^{2}=JE_{J}+\Delta E_{\Delta},$ (1)
where the spin variables $\sigma_{x}$ take on the values $-1,0$, or $+1$, $\langle xy\rangle$ indicates summation over nearest neighbors, and $J>0$ is the ferromagnetic exchange interaction. The parameter $\Delta$ is known as the crystal-field coupling and it controls the density of vacancies ($\sigma_{x}=0$). For $\Delta\rightarrow-\infty$, vacancies are suppressed and the model becomes equivalent to the simple Ising ferromagnet. We point out that the decomposition on the right-hand side of Eq. (1) into the bond-related and crystal-field-related energy contributions $E_{J}$ and $E_{\Delta}$, respectively, will be useful in the context of the numerical methods discussed below.
The phase boundary of the model in the crystal field–temperature ($\Delta,T$) plane separates the ferromagnetic from the paramagnetic phase. The ferromagnetic phase is characterized by an ordered alignment of ±$\pm 1$ spins. The paramagnetic phase, on the other hand, can be either a completely disordered arrangement at high temperature or a ±$\pm 1$-spin gas in a 00-spin dominated environment for low temperatures and high crystal fields. At high temperatures and low crystal fields, the ferromagnetic–paramagnetic transition is a continuous phase transition in the Ising universality class, whereas at low temperatures and high crystal fields, the transition is of first order [1, 2]. The model is thus a paradigmatic example of a system with a tricritical point $(\Delta_{\rm t},T_{\rm t})$ [3], where the two segments of the phase boundary meet. At zero temperature, it is clear that ferromagnetic order must prevail if its energy $zJ/2$ per spin (where $z$ is the coordination number) exceeds that of the penalty $\Delta$ for having all spins in the ±$\pm 1$ state. Hence the point $(\Delta_{0}=zJ/2,T=0)$ is on the phase boundary [2]. On the other hand, for zero crystal fields, the transition temperature is not exactly known, but it has been well-studied for a number of lattice geometries. For a picture of the phase diagram based on the data of the current work but also on some previous computations from Ref. [8], we refer the reader to Fig. 1. For brevity, we set $J=1$ and $J/k_{\mathrm{B}}=1$ as the unit of energy and temperature hereafter.
@@
Figure 1: Phase diagram of the two-dimensional spin-$1$ triangular Blume-Capel ferromagnet. (a) Transition points in the plane of the crystal-field coupling $\Delta$ and temperature $T$. The first-order transition line and the location of the tricritical point are determined from the field-mixing analysis using the Wang-Landau (WL) joint density of states. The critical points in the second-order transition regime are computed through multicanonical (MUCA) calculations in this work. For completeness, the earlier Wang-Landau simulation results from Ref. [8] are also presented. We note that the first-order transition points from the multicanonical calculations are omitted, as they are virtually identical to the Wang-Landau results. The complete list of the transition points is given in Table 1. (b) The phase diagram is replotted in the $\beta=1/T$ and $\mu=\Delta/T$ plane to illustrate the two scaling fields, $g$ and $\lambda$, from the field-mixing analysis described in Sec. II.3. (c) An enlargement of the phase boundary near the tricritical point is used to estimate the field-mixing parameter $r$ from its slope.
In this paper we are interested in the critical properties of the Blume-Capel model embedded in the triangular lattice, where $z=6$ and $\Delta_{0}=3$. For this particular case, Mahan and Girvin [35] were the first to apply position-space renormalization-group methods to solve the model. These authors estimated the critical frontier with the location of the tricritical point $(\Delta_{\rm t},T_{\rm t})=(2.686,1.493)$. Many years later, Du et al.[36] performed a further sophisticated analytical calculation using an expanded Bethe-Peierls approximation to find respectively the estimates $(\Delta_{\rm t},T_{\rm t})=(2.841,1.403)$ [37]. The refinement $(\Delta_{\rm t},T_{\rm t})=(2.925(8),1.025(10))$ was proposed in Ref. [8] via Monte Carlo simulations and an approximate scheme that takes advantage of the scaling behavior of the specific heat. As a historical note, we recall that in the 1980s and 1990s, the spin-$1$ Ising model on the triangular lattice was utilized as a $3$-state lattice-gas model, drawing inspiration from experiments on electrolyte adsorption on single-crystal (111) surfaces of FCC metals. While experimental interest primarily focused on the antiferromagnetic case, comprehensive phase diagrams for both ferromagnetic and antiferromagnetic models, highlighting topologically distinct ground states, can be found in Refs [38, 39]. For a more recent study of the triangular antiferromagnetic Blume-Capel model we refer the reader to Ref. [40].
@@ II Tricriticality
Figure 2: A comparison of the logarithm of the joint density of states, $\ln\Gamma(E_{J},E_{\Delta})$, for the triangular (a) and square-lattice (b) Blume-Capel models, as obtained through Wang-Landau simulations for a system with linear size $L=20$.
@@ II.1 Wang-Landau joint density of states
Figure 2 compares the landscape of $\Gamma(E_{J},E_{\Delta})$ for the triangular and square-lattice systems, revealing a spiky structure at high $E_{J}$ in the triangular-lattice case that is absent in the square-lattice system [20, 24]. This spiky structure in the triangular-lattice system is linked to configurations exhibiting frustration. Although the frustration is specific to the antiferromagnetic model, the structure in $\Gamma$ influences the Wang-Landau process regardless of the sign of $J$, as the algorithm must explore the entire energy space, which can slow down the sampling dynamics necessary to satisfy the flatness criterion. Here, Wang-Landau data of the joint density of states were obtained for linear sizes $L=\{8,9,10,12,14,15,16,18,20,21,22,24\}$. Five (two) samples of the density of states were prepared from independent simulations for $L<20$ ($L\geq 20)$. The deviations between different samples turned out invisibly small for the data analyzed in the following sections. In computational time cost, the largest system size $L=24$ of our current simulations in the triangular model is comparable to the previous calculations conducted on the square-lattice model with $L=48$ [24].
@@ II.3 Method of field mixing
where $s$ and $r$ are the system size-dependent field-mixing parameters, and $\beta_{\rm t}$ and $\mu_{\rm t}$ indicate the tricritical point. The direction of the scaling field $g$ is tangent to the phase-coexistence curve. The field $\lambda$ is across the phase boundary but is not necessarily orthogonal to $g$. The directions of $\lambda$ and $g$ are graphically described in Fig. 1(b), where the field-mixing parameters are associated with the angles as $r=-\tan\theta$ and $s=\tan\Psi$. The scaling operators conjugate to $\lambda$ and $g$ are
@@
respectively, where $\epsilon\equiv L^{-d}E_{J}$ and $n\equiv L^{-d}E_{\Delta}$ with dimension $d=2$ in our model, satisfying the relation $\langle\mathcal{O}_{x}\rangle=-L^{-d}\partial_{x}\ln\mathcal{Z}$. Note that $n$ represents the density of non-zero spin states, and the operator $\mathcal{Q}$ deviates from $n$ when $s\neq 0$. For example, at $T=0.985$, our estimates are $s\approx 0.028$ for $L=24$ and $s\approx 0.085$ for $L=48$, as we discuss in the analysis below. The probability distribution of the operator $\mathcal{Q}$ is a crucial tool in the field-mixing analysis, enabling the determination of the phase-coexistence curve and the identification of the tricritical point.
At a first-order transition point, the probability distribution $P(\mathcal{Q})$ conjugate to the field $\lambda$ in the direction across the phase boundary is double-peaked and symmetric. On the other hand, the distribution of the operator $\mathcal{E}$ conjugate to the field $g$ along the tangent of the phase boundary is bell-shaped. Although both follow the scaling Ansatz near the tricritical point, it was demonstrated that the symmetric double-peaked structure of $P(\mathcal{Q})$ works as a sensitive indicator of the phase coexistence when computed using the Wang-Landau joint density of states [24].
@@
Although the double-peaked structure is evident at temperatures both below and above the tricritical temperature $T_{\rm t}$, as illustrated for small systems in Fig. 3, the definite first- and second-order character of the transition can be undoubtedly distinguished by the finite-size behavior of $P(\tilde{\mathcal{Q}})$. At the first-order transition regime the peaks become sharper as $L$ increases (see also Figs. 9(a) and 9(b) below). In contrast, at the second-order transition, as shown in Fig. 3(d), the double-peaked structure becomes less pronounced and the peak height appears to decrease with increasing system size. Within the scope of the present calculations, we are unable to make a definitive prediction regarding whether the double-peaked structure persists in the thermodynamic limit. Addressing this question would require numerical data for very large systems, which are beyond the reach of our current computational resources.
@@ II.5 Tricritical exponents
The identification of the tricritical point in our field-mixing analysis prompts us to verify the proposed tricritical Ising exponents. Near the tricritical point, the probability distribution can be expressed as
@@
where $p^{\ast}_{L}$ is a universal scaling function, $m$ and $h$ denote the magnetization and magnetic field, respectively, and $a$’s are non-universal factors (for further details on the field-mixing method we refer the interested reader to Refs. [47, 46, 45, 48]). The relevant eigenvalue exponents for the tricritical Ising universality class in two dimensions are known to be $y_{t}=9/5$, $y_{g}=4/5$, and $y_{h}=77/40$ [51, 52, 53, 54].
At the tricritical point, the finite-size scaling Ansatz for the variable $\mathcal{Q}$ can be expressed more explicitly as
@@
where $\mathcal{P}_{o}$ is scale-invariant. To examine this scaling form, it is essential to have both mixing parameters, $r$ and $s$, to compute $\mathcal{Q}$. While we have obtained $s$ from the field-mixing analysis, the other parameter $r$ can be determined from the phase coexistence curve in the $\beta-\mu$ plane. In Fig. 1(c) we obtained $r=-3.2009(3)$ from the slope of the curve at the tricritical point. Figure 5(a) shows an excellent finite-size scaling collapse of $P_{L}(\mathcal{Q})$ with the given exponent $y_{t}=9/5$ at $T=0.985$, verifying the conjectured tricritical exponent in the triangular Blume-Capel ferromagnet.
@@ III.1 Multicanonical simulations
In our multicanonical simulations, we considered triangular lattices with periodic boundary conditions and linear sizes in the range $12\leq L\leq 128$. To thoroughly probe the characteristics of the transition on both sides of the phase boundary across the tricritical point, we performed simulations at various temperatures from $T=2.395$ in the second-order transition regime down to $T=0.75$ in the first-order regime. We note here that the multicanonical simulations become prohibitively demanding in the low-temperature regime, a fact that is also reflected in the statistical fluctuations of the large-$L$ data illustrated in Figs. 10 and 11 for $T=0.75$. Last but not least, for all fits performed throughout this paper, we restricted ourselves to data with $L\geq L_{\rm min}$, adopting the standard $\chi^{2}$ test for goodness of the fit. Specifically, we considered a fit as being acceptable only if $10\%<Q<90\%$, where $Q$ is the quality-of-fit parameter [68].
@@ III.2 Observables
At this point, having defined some of the main observables of interest, we call attention to a comparative check that highlights the numerical accuracy of the simulation methods implemented in this work. Figure 8 contrasts the specific heat and magnetic susceptibility curves of a $L=16$ triangular Blume-Capel system at the temperature $T=2.395$ obtained via three different protocols: (i) the Wang-Landau joint density of states outlined in Sec. II.1, (ii) the multicanonical scheme sketched here, and (iii) the well-known Metropolis algorithm [55]. Without doubt, Fig. 8 showcases a perfect matching among the three datasets.
As we are mostly interested in the first-order characteristics of the transition, following the prescription of Ref. [64], we constructed the probability density function of $E_{\Delta}$ per spin, $P(e_{\Delta})$, where $e_{\Delta}=E_{\Delta}/L^{2}$. It is well known that a double-peaked structure in the probability density function in finite systems is a precursor of the two delta functions expected in the thermodynamic limit for a first-order phase transition [49, 50]. Typical (normalized to unity) probability distributions $P(e_{\Delta})$ for various system sizes and temperatures in the first-order transition regime and at the proposed tricritical point are shown in Fig. 9.
A few comments are in order: (i) In the first-order transition regime, the barrier increases upon lowering the temperature, reaching the order of $\sim 10^{-24}$ for $T=0.75$ [panel (a)]. (ii) Note the strong enhancement of the barrier with increasing system size [panel (b)], exhibiting a major suppression of intermediate states, which is common for first-order phase transitions. (iii) At the proposed tricritical point, [panel (c)], the distributions still feature a double-peaked structure, yet with a much smaller barrier which seems to be independent of system size. A more illuminating discussion on the asymptotic behavior of these observables regarding the distributions of Fig. 9 is given below.
@@ III.3.1 First-order transition regime
where $P_{\rm max}$ and $P_{\rm min}$ are the maximum and local minimum of the distribution $P(e_{\Delta})$, respectively.
The resulting barrier connects a spin-00 dominated regime ($e_{\Delta}$ small) and a spin-±$\pm 1$ rich phase ($e_{\Delta}$ large). The corresponding surface tension $\Sigma(L)=\Delta F(L)/L$ is expected to scale as
@@
in two dimensions, possibly with higher-order corrections [75, 76, 77]. We point out here the recent numerical evidence of Ref. [78] that relate the surface tension $\Sigma$ to an interface between a spin-±$\pm 1$ phase separated by a surfactant layer of zeros, leading to a reduced effective surface tension [79, 80]. The scaling behavior of the surface tension is depicted in Fig. 10(a) for the complete set of temperatures considered during the multicanonical simulations and the solid lines are fits of the form (28). The retrieved $\Sigma_{\infty}$ values are plotted against the temperature difference from the tricritical point ($T_{\rm t}-T$) in Fig. 10(b), featuring an almost linear increase in $\Sigma_{\infty}$ as the temperature decreases towards the limit $T\rightarrow 0$. This result agrees nicely with the observations by Jung and Kim for the square-lattice Blume-Capel model; see Fig. 4(a) in Ref. [33]. Additionally, we verify that at the proposed tricritical temperature $T=0.985$, the surface tension vanishes to a limiting value of $\Sigma_{\infty}=3\times 10^{-6}\pm 2\times 10^{-5}$.
We now examine the system-size scaling behavior of the pseudo-transition field $\Delta_{L}^{\ast}$ obtained using the probability density function $P(e_{\Delta})$. We determine $\Delta_{L}^{\ast}$ by the value of the crystal field at which the two peaks of equal height appear in $P(e_{\Delta})$, as illustrated in Fig. 9. Figure 11 presents the behavior of $\Delta_{L}^{\ast}$ for selected temperatures in the first-order transition regime. Our data indicate an excellent fit to the expected first-order behavior of the scaling form (10) as demonstrated by the solid lines in Fig. 11. The extrapolated values of the transition field $\Delta_{\infty}^{\ast}$ obtained here from $P(e_{\Delta})$ confirm to a high numerical accuracy the transition points obtained from the field-mixing analysis in Sec. II.3.
Table 1 provides the comparison between the two different methods, showing that a deviation in $\Delta_{\infty}^{\ast}$ is about $10^{-4}$ or less.
@@
where $\{b_{C},b^{\prime}_{C},b_{\chi},b^{\prime}_{\chi}\}$ are non-universal amplitudes. This scaling Ansatz provides an excellent fit to our data for both observables, as shown in Fig. 12.
From the data fitting with $L_{\mathrm{min}}=24$, we obtain $x=2.000(2)$ and $2.000(4)$ for the specific heat and the magnetic susceptibility, respectively, which are in perfect agreement with the expected value $d=2$ of the first-order transition in this model. Note that due to the value of $x\approx 2$, the $1/L^{2}$ correction simply becomes an additive constant.